Use the normal approximation to the binomial to find the probability for n=10, p=0.5 and X >9.
The probability is...

Answer by Boreal(15235)

mean is 5
variance is np(1-p) or 2.5
sd is 1.58
use 9.5 as continuity correction factor
z>(9.5-5)/1.58=4.5/1.58=2.848
=0.0022 ANSWER
(the actual result is 0.0010.)